Please help me solve this word problem:
A computer store sells two different types of ink cartridges for a popular inkjet printer. The brand name cartridge sells for $28 and the generic cartridge sells for $12. Last week they sold 82 cartridges for that printer. If the total collected from the sale of these cartridges was $1,512, how many of each type of cartridge were sold?

Answer by checkley77(12569)

X+Y=82 OR X=82-Y
28X+12Y=1,512 REPLACE X BY (82-Y) IN THIS EQUATION & SOLVE FOR Y.
28(82-Y)+12Y=1,512
2,296-28Y+12Y=1,512
-16Y=1,512-2,296
-16Y=-784
Y=-784/-16
Y=49 ANS.
X+49=82
X=82-49
X=33 ANS.
PROOF:
28*33+12*49=1,512
924+588=1,512
1,512=1,512